OpenStudy (anonymous):

what is one solution of the equation(sin(X)+cos(X))^2=2? multiply choice !!!

OpenStudy (tkhunny):

Have you considered squaring the part inside the parentheses? Perhaps you have considered transforming first? \(\sin(x) + \cos(x) = \sqrt{2}\cdot \cos(x-\pi/4)\)

OpenStudy (anonymous):

i got \[\sin ^{2}x+2sinxcosx+\cos ^{2}x=2\]

OpenStudy (ranga):

\[\sin^2(x) + \cos^2(x) = 1; \quad 2\sin(x)\cos(x) = \sin(2x)\]

OpenStudy (anonymous):

\[1+2sinxcosx=2\]

OpenStudy (anonymous):